As part of your physics class, you calculated your automobile’s average fuel usage over a short interval of time and found it to be 28 MPG. If you drive an average of 120 miles per week and the cost of gasoline averages $2.75 per gallon, how much do you spend per year on gasoline?
1
Expert's answer
2020-09-16T10:12:41-0400

If you drive 120 miles per week then you spend "\dfrac{120}{28} = \dfrac{30}{7}" gallons of fuel per week. Then, during a year you will spend "\dfrac{30}{7}\times \dfrac{365}{7} =\dfrac{10950}{49}" gallons (there are "365\/7" weeks in a year). Then, the price will ge:

"p = \dfrac{10950}{49} \times \$2.75 \approx \$614.54"

Answer. $614.5.